OpenStudy (anonymous):

the components of vector are Ax = 3.9 and Ay = -4.0. What is the angle measured counterclockwise from x-axis to vector a?

OpenStudy (aaronq):

 \(tan\theta=\dfrac{4}{3.9}\rightarrow \theta=tan^{-1}(\dfrac{4}{3.9})=45.73^o\)

OpenStudy (anonymous):

Great thank you!